OpenStudy (anonymous):

I forgot how to do this, I'm in eight grade, help please. Solve the equation. 0.8=a/6.5

OpenStudy (anonymous):

a= 5.2

OpenStudy (anonymous):

Thanks =]

OpenStudy (phi):

multiply both sides by 6.5 \[ 6.5 \cdot 0.8 = \frac{a \cdot 6.5}{6.5} \] the idea is that 6.5 divided by itself is 1 so you get \[ 6.5 \cdot 0.8 = a \] you need to multiply out 6.5 times 0.8 to get the final answer

OpenStudy (anonymous):

Thanks for the explanation phi =)